Netflix is selling for $100 a share. A Netflix call option with one month until expiration and an exercise price of $105 sellPlease do 5B
Answer #1

1.

Using put call parity

Price of bond=S+P-C=100+6.94-2=104.94

2.

Effective annual interest rate is given as equal to=(105/104.94)^12-1=0.7%